Consider two dimensional square lattice site of side 0.5nm.At what electron

momentum values do the sides of the first Brillouin zone come? What is the

energy of free electron with this momentum.


1
Expert's answer
2022-03-29T12:29:14-0400

Answer

electron

momentum values

"P=\frac{h}{2a}\\=\frac{6.626*10^{-34}}{2*0.5*10^{-9}}\\=6.626*10^{-25}Kg.m/sec."